An electron is accelerated through a 100 V potential difference before perpendicularly entering a uniform magnetic field. Inside the field, the electron moves in circular trajectory of radius 4 cm, calculate the magnetic force that acts on the electron.


1
Expert's answer
2022-02-22T10:30:23-0500

"B=\frac{mv}{eR},"

"v=\sqrt{\frac{2eU}m},"

"B=\sqrt{\frac{2mU}{eR^2}}=5.2~mT."